How many pounds of candy selling for $1.75 a pound should be added to three pounds of candy selling for $4.25 a pound to create a blend, which will sell for $3.25 a pound?
Answer by Edwin McCravy(20055)

How many pounds of candy selling for $1.75 a pound should be added to three pounds of candy selling for $4.25 a pound to create a blend, which will sell for $3.25 a pound?
Let x = the number of pounds of the cheaper candy to be mixed
Let y = the total number of pounds of candy blend.

First you have the AMOUNT OF CANDY equation:

+= or x + 3 = y  

And second, you have the AMOUNT OF MONEY equation:

+= or (x)(1.75) + (3)(4.25) = (y)(3.25) or 1.75x + 12.75 = 3.25y

So you have this system of equations:

x + 3 = y
1.75x + 12.75 = 3.25y
 

Solve that system and get x = 2 and y = 5.

Answer: Add 2 pounds of the more expensive candy to end up with 
5 pounds of candy blend. 

Edwin